Project

General

Profile

Actions

20160202 » History » Revision 5

« Previous | Revision 5/7 (diff) | Next »
Boyd, Steven, 03 February 2016 11:03


Commissioning and Ops Scheduling meeting : 2/2/2016

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Action Item : John will look at the effect of air on the MC truth scattering angle
    • Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
      • Action Item : DC to work with Pierrick to ensure this
    • Action Item : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
    • Action Item : Paolo, Mike and Vicky will investigate the Linde fridge PC instability issue
    • Action Item : Paolo/Pierrick will think of a way to monitor the Linde fridge PC
    • Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
      and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
  • Current schedule
    Feb 1 - 5 : Daresbury Electrical work ( 10 A test )
    Feb 2 : DS coldbox cool down
    Feb 4 : DS magnet cool down
  • Updates and future schedule
    • DS Cooldown schedule
    • Realignment work and schedule
    • Cooling water discussion
    • Controls and Monitoring : Network reconfiguration
    • Tracker patch panel leak work
    • Electrical work schedule
    • Data-taking schedule
      • User cycle begins Feb 16th
      • Start-up activities : Feb 16 - 21
      • Data-taking starts : Feb 23

Minutes

Present: Steve B, Chris R, Paul H, Andy N, Vicky B, Mike C, Ryan B, Paul K, Paolo F, Josef F, Henry N, Mark T, Melissa U, Victoria B, John N

*Action Items
  • Action Item : John will look at the effect of air on the MC truth scattering angle
    • John presented plots of MC scattering angle in air and vacuum. They contain some features that we don't understand. John will continue looking into these.
  • Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
    • Action Item : DC to work with Pierrick to ensure this
    • Action Item stands
  • Action Item : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
    • Email sent but no reply. Steve will try again
  • Action Item : Paolo, Mike and Vicky will investigate the Linde fridge PC instability issue
    • The IOC needed to be restarted. This fixed the problem
  • Action Item : Paolo/Pierrick will think of a way to monitor the Linde fridge PC
    • Paolo has added a monitoring process to Nagios to ensure that the IOC on this PC is always alive
  • Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
    and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
    • Action Item postponed. However we do need to do this in the near future. Colin and Trevor spoke informally about priorities and Trevor looked through the schedule. The FC requires about 2 weeks of work. Once it's online some work becomes near-impossible. Trevor is focussing on the FC now. DL are also helping out on the specs of the dump resistors. Work being done on the SS is conditional on the results Steve gets from 10A test. Pierrick reported that the 10A tests are finishing up now. There is no real conclusion on this test. Did find that a small resistance in parallel to the magnet caused overshoots and undershoots so not properly damped. If issues can't be solved we may have to change where the resistors go. Next week most of the PS modification hardware is being sent to RAL so most of the work is done before March is done.
  • Issues of chain of authority during "mixed working" i.e. work in rack room at the same time as data-taking needs to be addressed.
  • We all agreed that the DC is in overall charge, even during data-taking. However, we need think about the key dance and what needs to
    be done to ensure clarity in the line of authority.
  • Action Item : Steve/DC/Andy to think about organisation during mixed work sessions
  • Feb 3 DS coldbox cool down
  • Feb 3 DS magnet cool down
  • DS cooldown about to start
  • Realignment work: Drawing for modifications to the bellows are ready to go out.
    Colin doesn't think we can the bellows ready for the data-taking time. Most sensible thing to do is to do as much work as possible but
    not move the solenoids now. THis will enable us to take data, before doing the realignment work during the next shutdown period after
    user cycle 2015/04. John thinks it's the right way to go. Victoria says that we are not running with fields so the alignment is not
    a priority.
    Ken : will we have He in the tracker volume.
    Colin : we need to know what the leak rate is in both trackers and we should look at them before data-taking.
    Next Ops meeting : Leak checking plan.
    Action Item : Colin/Me talk to Melissa about Helium leak fix
  • Cooling water discussion postponed
  • Realignment of PID detectors has been done. Need a survey. Should get the EMR in position. * Critical path : what is going on with the tracker. * Superconducting permit and PRY discussion :
    Action Item : Andy/Colin/Ken/DC
  • Patch panel work flipped to tracker group
  • Controls and Monitoring.
    All IOCs configured with static IPs. Still need to deploy list of machine/IP number. Each IOC needs to be tested. Network lockout
    code is being written. Will be done by the next meeting.
  • Electrical work in-hand. Electrical test being continued.

Henry : Test of TOF0, Cerenkov in case magnet flushing had an effect.

AOB:
Henry: reminds me to see Debbie about phone. Paul K will look at this.
Mike: is there any intention to power the FC for the forthcoming User Cycle. When should we put the FC on the frame.
Ken : Work plan for the Daresbury team. Action Item : Schedule a meeting between Colin/Trevor/Myself/Pierrick.
Henry : Move to R1. Advance party has gone. Move is not next week, but may be the week after. Colin will set a date for
the move.
Pierrick : reminder to Mark that they will look at the SSU Helium.

Updated by Boyd, Steven about 8 years ago · 5 revisions